OpenStudy (cicilybailey14):

Fan+Medal!! For the data in the table, tell whether y varies directly with x. If it does, write an equation for the direct variation. x y -2 1 3 6 8 11 No it does not. Yes; y = x + 3 Yes; y = -1/2x Yes ; y = -2x

OpenStudy (anonymous):

SLOP BEETWEN 2 POINTS: Y-Y0/X-X0==> 6-11/3-8 SO SLOP=5/5=1 SO WE HAVE: Y=ax+b so y=x+3